### v1

# Judging the Judges: Can Large Vision-Language Models Fairly Evaluate Chart Comprehension and Reasoning?

Abstract
Charts are ubiquitous as they help people understand and reason with data. Recently, various downstream tasks, such as chart question answering, chart captioning, etc. have emerged. Large Vision-Language Models (LVLMs) show promise in tackling these tasks, but their qualitative evaluation is costly and time-consuming, limiting real-world deployment. While using LVLMs as judges to assess chart comprehension capabilities of other LVLMs could streamline evaluation processes, challenges like proprietary datasets, restricted access to powerful models, and evaluation costs hinder their adoption in industrial settings. To this end, we present a comprehensive evaluation of 13 open-source LVLMs as judges for diverse chart comprehension and reasoning tasks. We design both pairwise and pointwise evaluation tasks covering criteria like factual correctness, informativeness, and relevancy. Additionally, we analyze LVLM judges based on format adherence, positional consistency, length bias, and instruction-following. We focus on cost-effective LVLMs ($\leq 9$B parameters) suitable for both research and commercial use, following a standardized evaluation protocol and rubric to measure the LVLM judge accuracy. Experimental results reveal notable variability: while some open LVLM judges achieve GPT-4-level evaluation performance (about 80% agreement with GPT-4 judgments), others struggle (below  10% agreement). Our findings highlight that state-of-the-art open-source LVLMs can serve as cost-effective automatic evaluators for chart-related tasks, though biases such as positional preference and length bias persist.
Judging the Judges: Can Large Vision-Language Models Fairly Evaluate Chart Comprehension and Reasoning?
Md Tahmid Rahman Laskar‡,thanks:   Contact Emails: {tahmedge,enamulh,jhuang}@yorku.ca, Mohammed Saidul Islam‡,thanks:   Equal Contributions., Ridwan Mahbub‡,22footnotemark: 2, Ahmed Masry, Mizanur Rahman, Md Amran Hossen Bhuiyan, Mir Tafseer Nayeem§, Shafiq Joty, Enamul Hoque‡,11footnotemark: 1, Jimmy Xiangji Huang‡,11footnotemark: 1 York University, §University of Alberta, Salesforce AI Research
Figure 1: An example evaluation of Gemini-1.0-Pro model generated L2/L3 caption in the VisText dataset by an LVLM judge: the LLaVA-Critic-7B model.
1 Introduction
Understanding data visualizations—such as bar and line charts—requires multimodal reasoning, as it involves integrating visual encodings with textual and contextual information Hoque and Islam (2024). Recent research has introduced various tasks (e.g., chart question answering, chart captioning, fact-checking with charts, etc.) to facilitate chart-based reasoning via natural language. These tasks demand the understanding of both the chart’s visual content (data values, trends, visual encodings) and accompanying text or instructions.
Large Language Models (LLMs) have revolutionized NLP and vision-language tasks Zhao et al. (2023), with growing interest in their use for chart comprehension and reasoning due to their strong multimodal capabilities. This progress can have a substantial impact on real-world industrial applications, where extracting insights from charts and graphs can drive critical business decisions Obeid and Hoque (2020); Masry et al. (2023); Meng et al. (2024). However, evaluating LLM performance in chart understanding presents notable challenges Islam et al. (2024). For instance, traditional text-based metrics like BLEU fail to capture the quality of open-ended explanatory answers and also require human-annotated references. While human evaluation can address this problem, it is time-consuming and resource-intensive.
To address this, recent studies have proposed using LLMs themselves as automatic evaluators or judges Gu et al. (2024); Li et al. (2024b). By employing LLMs to evaluate the chart comprehension abilities of other models (see Figure 1 for an example), the evaluation process can be streamlined, making the process more efficient and reproducible without human intervention. While this method accelerates development and reduces dependency on human annotations, its real-world adoption is hindered by privacy and scalability constraints. For example, organizations may be unwilling to share proprietary data with closed-source models from OpenAI, Google, or Anthropic. While closed-source models demonstrate impressive judging capabilities, their compatible open-source models are often large in size (e.g., 70B to 400B parameters). This requires high computing resources and usage costs. Therefore hinders real-world utilization.
To this end, this paper aims to investigate whether open-source smaller LVLMs (e.g., less than 10B parameters) can effectively evaluate answers about charts – assessing correctness, relevance, and other qualities – similarly to a human or a powerful LLM like GPT-4 OpenAI et al. (2023)? For this purpose, we conduct one of the first comprehensive evaluations of open-source LVLMs as evaluators on various chart benchmarks, consisting of diverse tasks like chart captioning and question answering. We focus on open-source, smaller VLMs (up to 10B parameters) to simulate realistic deployment scenarios where cost-effective or private models are preferred over large closed models. By benchmarking these models against high-quality reference judgments generated by closed-source LLMs like GPT-4 or 70B open-source LLM-Judge like LLaVA-Critic Xiong et al. (2024), we aim to uncover to what extent current open models can serve as reliable judges, and when they fail.
Our major contributions to this paper are:
(1) We establish an evaluation framework for chart comprehension using “LVLM-as-a-Judge”, with clear rubrics for pairwise and pointwise assessments over 100K judgments generated by GPT-4o and LLaVA-Critic-70B. Additionally, we introduce a new benchmark to assess the instruction-following abilities of LVLMs in Chart related tasks.
(2) We evaluate a wide range of open-source multimodal LLMs as judges – 13 models ranging from 2B to 9B parameters – and analyze their performance against LLM-annotated (GPT-4 and LLaVA-Critic) and human-annotated reference judgments, across diverse chart benchmarks (OpenCQA and VisText) on answers generated by different LLMs to create challenging evaluation scenarios.
(3) We provide an in-depth analysis of the judges’ strengths and weaknesses, revealing issues like position bias and length bias, and discuss which models achieve substantially higher agreement with reference judgments, and which ones fail.
In addition, our code, judgment data, and our proposed instruction-following evaluation benchmark is released here: https://github.com/tahmedge/chart_lvlm_judge.
2 Related Work
Earlier efforts in chart question answering include synthetic datasets like FigureQA Kahou et al. (2017) and DVQA Kafle et al. (2018), which generated templated QA pairs for simple charts but lacked real-world complexity. ChartQA Masry et al. (2022) addressed this gap with real-world charts and more complex questions, while OpenCQA Kantharaj et al. (2022) pushed further with open-ended, explanatory queries. Meanwhile, chart captioning has emerged as another avenue for summarizing chart content Shankar et al. (2022); Rahman et al. (2023); Tang et al. (2023). Together, these datasets highlight the growing complexity of chart-based reasoning tasks and the need for more robust evaluation methods.
While the rise of multimodal LLMs offers potential for chart-related tasks, general vision-language models often struggle with chart-specific elements like axis text and precise data points Islam et al. (2024). To address this, specialized models such as ChartLLaMA Han et al. (2023), ChartInstruct Masry et al. (2024), ChartGemma Masry et al. (2025), and TinyChart (Zhang et al., 2024) have been developed, showing strong performance. However, evaluating these models is challenging, as many still depend on time-consuming human assessments for open-ended responses.
While using LLMs to evaluate other LLMs has gained a lot of attention, early efforts focused primarily on text-only tasks like summarization Li et al. (2024b); Zheng et al. (2023). For multimodal tasks, models such as Prometheus-VL Lee et al. (2024) and LLaVA-Critic Xiong et al. (2024) introduced smaller open-source vision-language models (as small as 7B) fine-tuned to serve as general-purpose multimodal evaluators. Our work aligns with this direction, leveraging LVLMs as judges. Although concurrent studies explore similar capabilities Chen et al. (2024), they report that early LVLMs like LLaVA-1.5 struggle with text-rich visuals such as charts and diagrams Lee et al. (2024). Addressing the gap in evaluating recent LVLMs on chart-specific tasks, we present the first systematic study of state-of-the-art open-source LVLMs as judges across diverse chart comprehension and reasoning benchmarks.
Figure 2: An overview of our evaluation methodology.
3 Methodology
Given a chart and model generated response(s), we construct the prompt (see Appendix A.2 for some sample prompts) depending on the evaluation rubric. Following the prior work Lee et al. (2024), we ask the LVLM judge to provide their answer along with an explanation, since adding an explanation during assessments ensured better judgment performance in early work. Below, we describe our evaluation method (also see Figure 2).
3.1 Evaluation Rubric Design
Following the prior work on LLM evaluation Chen et al. (2024); Lee et al. (2024); Xiong et al. (2024), we define clear rubrics for the judges:
(i) Based on Evaluation Type:
- Pairwise: The judge must select the better answer between two given responses (e.g., Claude vs Gemini) about the chart for the given instruction.
- Pointwise: The judge must rate a single answer to the chart query on a Likert scale from 1 (very poor) to 5 (excellent).
(ii) Based on Reference Type:
- With Reference: The judge is also given the ground-truth answer or summary as a reference, and instructed to choose the response that better matches the reference as well as the chart context.
- Without Reference: The judge only sees the model response(s) and the chart image and must decide based on its own judgment.
(iii) Based on Evaluation Criteria:
- Factual Correctness: Focuses only on the factual accuracy of the response.
- Informativeness: Focuses on the amount of useful information in the response.
- Relevance: Focuses on measuring the relevancy of the response.
- Multidimensional Evaluation: Considers overall response quality based on factual correctness, informativeness, conciseness, and relevance.
(iv) Based on Evaluation Metrics:
- Judgment Accuracy: The percentage of instances where the answer picked by the judge same as the gold. It is relevant to the pairwise case.
- Error Distance: The average absolute difference between the judge’s 1–5 rating and the reference’s rating. It is relevant to the pointwise case.
- Positional Bias Metric: In the pairwise case, we swapped the order of answers and checked if the judge’s decision changed.
- Length Bias Metric: Checked if the judge’s wrong choice correlated with the answer length.
- Instruction Following Evaluation Accuracy: We analyzed whether the LVLM judge can effectively evaluate the instruction following capability of other LVLMs.
- Format Adherence Accuracy: This metric measures whether the judge’s output followed the required JSON format.
3.2 Evaluation Data Construction
OpenCQA Kantharaj et al. (2022): This is an open-ended question-answering dataset on real charts. Each data point includes a chart and a question and expects an explanatory answer. We use its test set containing 1.1k QA instances.
VisText Tang et al. (2023): This is a chart captioning dataset with 12,441 charts, each paired with two types of captions: synthetic Level 1 (L1) captions that describe the chart’s structural elements—such as chart type, title, axis labels, and scales—and human-generated Level 2/Level 3 (L2/L3) captions that provide insights into key statistics, trends, and patterns within the data. We use both L1 and L2/L3 captions with 1.2K test instances for each type.
For OpenCQA and VisText, we use the outputs generated by Islam et al. (2024) using Gemini-1.0-Pro Team et al. (2023) and Claude-3-Haiku Anthropic (2024) and compute the judgment scores using GPT-4o OpenAI et al. (2023) and LLaVA-Critic-70B Xiong et al. (2024) models and use as the judgment reference for diverse scenarios, as demonstrated in the previous section. This results in about 100K judgment data generated by GPT-4o and LLaVA-Critic-70B. We select these two models due to their impressive performance as a multimodal LLM-Judge Xiong et al. (2024).
Chart-Instruct-Eval: We find that there are no datasets currently available in the chart domain that can assess the instruction-following capabilities of LVLMs. Therefore, we construct an instruction-following dataset (denoted as Chart-Instruct-Eval) to evaluate whether LVLM judges can evaluate the instruction-following capabilities of different models in chart-related tasks. For the dataset construction, we sample 400 charts from the ChartGemma Masry et al. (2025) dataset. However, the original input instructions in the ChartGemma dataset lacked sufficient details. Hence, we could not use it for the instruction following purpose. Therefore, for each sample, we first create a detailed instruction containing specific requirements for the LLM response in terms of formatting, length, and structure to ensure instruction following. Then we manually prepare one good and one bad response corresponding to the instruction. Both responses convey similar content, but the good response fully adheres to all provided instructions, whereas the bad response disregards them. Finally, we assess the LLM judges whether they can reliably evaluate which response properly follows the instructions.
3.3 LVLM Judges
We evaluate 13 different open-source multimodal LLMs [^1] We did not use the Prometheus-VL-7B Lee et al. (2024) model since it requires a specific input format, making our prompts incompatible. as candidate judges, focusing on relatively smaller, publicly available models (2B–10B parameters). These include: (i) XGen-MM-Phi-3-3.8B Xue et al. (2024)a multimodal model (3.8B) developed by Salesforce, (ii) MiniCPM-V-2.6-7B Yao et al. (2024)a 7B vision-language model by OpenBMB, (iii) Ph-3.5-3.8B-Vision-Instruct Abdin et al. (2024)a smaller vision model from Microsoft, (iv) Qwen2-VL-2B - Alibaba’s Qwen Wang et al. (2024) multimodal model with just 2B parameters, (v) Qwen2-VL-7BThe 7B version of the multimodal Qwen model, (vi) PaliGemma-3B Beyer et al. (2024)Google’s multimodal open-source model, (vii) ChartGemma Masry et al. (2025)a chart-specialized model based on PaliGemma that is fine-tuned on chart tasks, (viii) Idefics-9B-Instruct [^2] HuggingFaceM4/idefics-9b-instructan open multimodal model known for image understanding, (ix) InternLM-XComposer-7B Dong et al. (2024)a 7B vision model with composition abilities, (x) LLaVA-v1.6-Mistral-7BA multimodal LVLM based on the LLaVA Li et al. (2024a) architecture that also utilizes a 7B Mistral Jiang et al. (2023) as the backbone, (xi) LLaVA-Critic-7Ba specialized evaluator model based on LLaVA and Qwen, (xii) mPLUG-Owl-3-7B Ye et al. (2023)a 7B multimodal model from Alibaba, (xiii) Janus-Pro-7B Chen et al. (2025) - an open-source LVLM developed by Deepseek. For more information about model selection, see Appendix A.1.
4 Experiments
Pairwise Evaluation: Judgment Accuracy (Higher is Better) Pointwise Evaluation: Error Distance (Lower is Better)
Model OpenCQA VisText L1 VisText L2/L3 OpenCQA VisText L1 VisText L2/L3
GPT-4o LC-70B Avg. GPT-4o LC-70B Avg. GPT-4o LC-70B Avg. GPT-4o LC-70B Avg. GPT-4o LC-70B Avg. GPT-4o LC-70B Avg.
Qwen2-VL-2B-Instruct 51.6 56.3 54.0 28.5 25.9 27.2 2.5 3.4 3.0 1.0 0.9 1.0 2.0 2.1 2.1 1.1 0.6 0.9
PaliGemma-3B 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 5.0 5.0 5.0 5.0 5.0 5.0 5.0 5.0 5.0
ChartGemma-3B 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 5.0 5.0 5.0 5.0 5.0 5.0 5.0 5.0 5.0
Phi-3.5-Vision-3.8B-Instruct 49.5 51.9 50.7 72.5 66.4 69.5 43.6 55.3 49.5 0.7 0.8 0.8 1.4 1.6 1.5 1.1 0.9 1.0
XGen-MM-Phi3-3.8B-Instruct 67.6 75.5 71.6 78.5 72.2 75.4 63.9 77.4 70.7 1.0 0.7 0.9 1.3 1.5 1.4 1.0 0.4 0.7
Janus-Pro-7B 46.6 48.7 47.7 48.6 45.6 47.1 52.6 57.0 54.8 1.0 0.7 0.9 1.0 1.2 1.1 1.0 0.4 0.7
Qwen2-VL-7B-Instruct 67.3 66.4 66.9 64.0 51.1 57.6 69.6 70.3 70.0 0.8 0.6 0.7 0.6 0.5 0.6 0.9 0.5 0.7
InternLM-Xcomposer2d5-7B 64.8 64.1 64.5 76.8 67.2 72.0 69.7 81.4 75.6 0.8 0.9 0.9 0.8 0.9 0.9 0.9 0.4 0.7
LLaVA-Next-v1.6-Mistral-7B 72.0 79.8 75.9 78.4 71.7 75.1 66.7 83.4 75.1 0.9 0.6 0.8 1.3 1.5 1.4 1.1 0.6 0.9
LLaVA-Critic-7B 75.1 83.8 79.5 82.8 75.3 79.1 69.0 85.1 77.1 0.5 0.4 0.5 0.5 0.4 0.5 0.8 0.4 0.6
mPLUG-Owl3-7B 60.8 59.4 60.1 72.2 64.0 68.1 46.1 39.2 42.7 0.8 0.6 0.7 1.0 1.0 1.0 0.9 0.4 0.7
MiniCPM-V-2.6-8B 64.3 68.6 66.5 49.2 42.9 46.1 44.8 39.1 42.0 1.0 0.8 0.9 1.3 1.3 1.3 1.7 1.5 1.6
Idefics-9B-Instruct 20.4 20.1 20.3 22.0 19.7 20.9 24.1 24.4 24.3 3.3 3.2 3.3 4.8 4.8 4.8 3.1 2.8 3.0
Table 1: Model performance based on average pointwise and pairwise scores across all reference types, as well as evaluation criteria (e.g., factual correctness, informativeness, etc.) in comparison to GPT-4o and LLaVA-Critic-70B (LC-70B) annotations (corresponding average score is also added). Bold values denote the best score in each case. Color coding for comparison: open-source models below 7B parameters, between 7-10B parameters.
In this section, we present the experimental results based on evaluating 13 LVLMs as judges across OpenCQA, VisText, and our proposed Chart-Instruct-Eval. The evaluation considers both pairwise and pointwise assessments, focusing on factual correctness, informativeness, relevance, positional bias, length bias, and instruction-following accuracy. We parse the LVLM-judge predicted judgments from their corresponding responses using a parsing script Laskar et al. (2023, 2024a, 2024b). Note that we ran all our experiments using 1 A100 GPU with all the decoding parameters being set to the default values in HuggingFace Wolf et al. (2020). Below, we demonstrate our findings:
4.1 Pairwise Evaluation Results
The pairwise evaluation measures how often the LVLM judges agree with GPT-4 or LLaVA-Critic-70B to select the better response in comparative assessments. We summarize the result in Table 1.
i. Top-performing models: LLaVA-Critic-7B achieved the highest agreement with reference judgments (above 75% average accuracy in each dataset). Another similar-sized (7B) LLM, the LLaVA-Next-v1.6-Mistral-7B model also performed competitively by exceeding 70% accuracy across each dataset. Interestingly, the XGen-MM model with just 3.8B parameters also achieved more than 70% accuracy, making it a very suitable judge in resource-constrained scenarios.
ii. Lower-performing models: Surprisingly, PaliGemma-3B and ChartGemma-3B achieved 0% agreement, indicating a poor alignment with reference judgments. Moreover, while the Qwen-2B model achieves decent performance in OpenCQA (above 50% accuracy), it achieves quite poor performance in VisText, especially in the L2/L3 scenario (below 10% accuracy). More surprisingly, the largest LVLM in our evaluation, the Idefics-9B-Instruct model achieves average accuracy below 25% in all datasets, highlighting its ineffectiveness as a judge. Our manual analysis revealed that these models failed due to not following instructions properly while also generating the response in the wrong format (improper JSON outputs). For PaliGemma, since it is not an instruction-tuned model, its poor performance could be related to the lack of understanding of instructions. The poor performance behind ChartGemma could be related to its training data lacking instructions related to judging tasks, therefore leading to poor generalization. We demonstrate some error examples of these LVLMs in Appendix A.3.
4.2 Pointwise Evaluation Results
This primarily measures the error distance between the ratings of the LVLM judge and the reference (GPT-4/LLaVA-Critic-70B) on a 1–5 Likert scale.
i. Top-performing models: Similar to the pairwise scenario, we find from Table 1 that LLaVA-Critic-7B again achieved the best performance in the pointwise scenario, achieving an error distance around 0.5. Other models like InternLM-Xcomposer-7B and Qwen2-VL-7B-Instruct that achieve quite good performance in pairwise scenarios, also demonstrate less error distance in pointwise scenarios (error distance below 1.0). Some other top-performing models in the pointwise scenario are LLaVA-Next-v1.5-Mistral-7B and MiniCPM-V-2.6-8B, which also achieve an error distance below 1.0 in 2 out of the 3 datasets.
ii. Lower-performing models: Similar to the pairwise scenario, PaliGemma-3B and ChartGemma-3B again produced irrelevant outputs resulting in the highest error distances (5.0). Moreover, despite being the largest model in our evaluation, the Idefics-9B-Instruct model performs quite poorly with very high error distance (on average, above 3).
4.3 Instruction and Format Adherence
We also assess the LVLM judges on their ability to maintain a standardized response format and whether they can evaluate the instruction following capabilities of other models. Based on the results in Table 2, we find that all 7B models achieve more than 90% format following capability. Smaller LVLMs like Qwen-2B and Phi-3.8B also achieve around 80% format adherence. In terms of instruction following capability evaluation, mPLUG-Owl3-7B (93.5%) and Qwen2-VL-7B-Instruct (87.0%) achieve the top two results. Surprisingly, the LLaVA-Critic-7B model achieves only 45.5% accuracy in this task. This may indicate that the training data of the LLaVA-Critic-7B model may not contain such data, leading to a quite poor generalization in this dataset. Moreover, PaliGemma-3B and ChartGemma-3B failed to follow the format requirements at all (also unable to evaluate instruction following capability). Finally, the Idefics-9B-Instruct model, even with 9B parameters achieves poor instruction and format following accuracy.
4.4 Bias Analysis
To assess potential biases in LVLM judges, we analyzed position bias (whether the order of the responses affects judgments) and length bias (whether longer responses are favored). Based on the result presented in Table 3, we find that the Qwen2-VL-7B-Instruct model exhibited the lowest positional bias and length bias. On the contrary, the LLaVA-Next-v1.6-Mistral-7B model showed very high bias in both scenarios, suggesting susceptibility to judge responses based on variations in the position of the responses as well as the length. Surprisingly, the LLaVA-Critic-7B model, which is the best-performing model in terms of judgment accuracy and error distance, demonstrates the highest length bias across all models, indicating a tendency to favor longer answers. We provide an example of the position bias in Figure 5, and an example of the length bias in Figure 6.
Model Instruction Following Format Adherence
Qwen2-VL-2B-Instruct 13.5 78.9
PaliGemma-3B 0.0 0.0
ChartGemma-3B 0.0 0.0
Phi-3.5-Vision-3.8B-Instruct 49.0 83.3
XGen-MM-Phi3-3.8B-Instruct 72.5 97.6
Janus-Pro-7B 73.0 96.7
Qwen2-VL-7B-Instruct 87.0 98.6
InternLM-Xcomposer2d5-7B 54.0 95.9
LLaVA- Next-v1.6-Mistral-7B 27.0 98.9
LLaVA-Critic-7B 45.5 99.7
mPLUG-Owl3-7B 93.5 98.9
MiniCPM-V-2.6-8B 54.5 90.3
Idefics-9B-Instruct 20.5 35.0
Table 2: Accuracy in terms of Instruction Following Evaluation (evaluated on Chart-Instruct-Eval) and Format Adherence (based on average across all datasets).
Model Length Bias Position Bias
Qwen2-VL-2B-Instruct 55.1 71.9
Phi-3.5-Vision-3.8B-Instruct 69.8 59.6
XGen-MM-Phi3-3.8B-Instruct 64.3 79.2
Janus-Pro-7B 27.2 50.6
Qwen2-VL-7B-Instruct 21.5 35.8
InternLM-Xcomposer2d5-7B 24.5 35.9
mPLUG-Owl3-7B 21.9 42.5
LLaVA-Next-v1.6-Mistral-7B 71.8 77.0
LLaVA-Critic-7B 76.4 39.6
MiniCPM-V-2.6-8B 37.4 45.5
Table 3: Length Bias and Position Bias for different models (results based on average across all datasets). Here, Lower values are better. Models achieving format following accuracy above 50% are only evaluated.
4.5 Human Evaluation
In this section, we conduct a human evaluation of the GPT-4o and the LLaVA-Critic-70B models which we used as the reference judge to evaluate the smaller open-source LVLMs. For this purpose, we randomly sample 100 responses generated by Islam et al. (2024) for the Claude-3-Haiku and the Gemini-1-Pro models in OpenCQA and VisText datasets. Then, we ask two human annotators having expertise in NLP and Computer Vision to rate these responses based on our evaluation criteria (e.g., informativeness, relevance, etc.) with references provided for 50% of the data and without any references for rest of the data.
Based on our human evaluation, we find that both annotators’ judgments highly correlate with GPT-4o and LLaVA-Critic-70B, with an error distance below 1.0. Interestingly, we find that both annotators have a higher correlation with the open-source LLaVA-Critic-70B model (average error distance with LLaVA-Critic-70B: 0.81, and with GPT-4o: 0.93). Therefore, in real-world industrial scenarios where human annotation is costly and closed-source LLMs are not preferred due to privacy concerns in proprietary datasets, the open-source LLaVA-Critic-70B model could be a good alternative for data annotation.
4.6 Ablation Studies
(i) Effect of Reference Type:
In this section, we compare the performance variation of different LVLMs in reference-based and reference-free scenarios (see Table 4). LVLMs that achieve more than 20% pairwise judgment accuracy in OpenCQA are selected for the analysis. While we find that different LVLMs have a slight change in performance with the presence and absence of references, the performance difference between them based on a paired t-test is not statistically significant ($p>0.05$). This demonstrates that the open-source LVLMs are robust in both reference-based and reference-free evaluation.
(ii) Effect of Evaluation Criteria:
In Table 5, we analyze the performance differences among various LVLMs with an error distance below $2.5$ in VisText (L1) across multiple evaluation metrics: (i) informativeness, (ii) relevance, and (iii) factual correctness. While we observe slight performance variations based on the evaluation criteria, the paired t-test demonstrates that these differences are not statistically significant ($p>0.05$), indicating robust performance across various evaluation measures.
Model With Reference Without Reference
Qwen2-VL-2B-Instruct 47.4 55.7
Phi-3.5-Vision-3.8B-Instruct 51.6 47.3
XGen-MM-Phi3-3.8B-Instruct 66.8 68.4
Janus-Pro-7B 45.9 47.3
Qwen2-VL-7B-Instruct 66.7 67.8
InternLM-Xcomposer2d5-7B 62.1 67.5
LLaVA-Next-v1.6-Mistral-7B 71.0 73.0
LLaVA-Critic-7B 74.9 75.3
mPLUG-Owl3-7B 63.5 58.2
MiniCPM-V-2.6 63.2 65.4
Idefics-9B-Instruct 16.6 24.2
Table 4: Judgment Accuracy in comparison to GPT-4o in OpenCQA based on Reference-based (with reference) and Reference-free (without reference) evaluation.
Model Factual Correctness Informativeness Relevancy
Qwen2-VL-2B-Instruct 2.6 1.6 2.0
Phi-3.5-Vision-3.8B-Instruct 1.6 1.3 1.4
XGen-MM-Phi3-3.8B-instruct-r-v1 1.7 1.4 1.6
Janus-Pro-7B 1.4 1.0 1.3
Qwen2-VL-7B-Instruct 0.7 0.4 0.5
InternLM-Xcomposer2d5-7B 1.0 0.8 0.9
LLaVA-Next-v1.6-Mistral-7B 1.7 1.4 1.5
LLaVA-Critic-7B 0.6 0.3 0.4
mPLUG-Owl3-7B 1.1 0.9 1.1
MiniCPM-V-2.6 1.7 1.3 0.9
Table 5: Average Error Distance (compared with LLaVA-70B-Critic) in VisText (L1) for different LVLMs based on various Evaluation Types. Here, lower values indicate better performance.
5 Conclusion and Future Work
In this paper, we conducted a comprehensive evaluation of open-source LVLMs as automatic judges for chart comprehension and reasoning tasks. Our analyses revealed that while some open-source LVLMs (e.g., 7B models like LLaVA-Critic, Qwen2-VL, InternLM, and LLaVA-Next) can achieve judgment accuracy (with lower error rates) which is comparable to state-of-the-art closed-source models like GPT-4 or larger open-source models like LLaVA-Critic-70B; other models, such as ChartGemma and PaliGemma, struggle significantly, highlighting variability in their reliability. Despite the promising results of various models, issues like bias and lack of instruction following capability still persist. Therefore, future work should focus on mitigating biases, improving instruction following evaluation capability, alongside ensuring consistency across diverse evaluation criteria by developing a multimodal LLM judge tailored for chart model evaluation.
Acknowledgments
This research was supported by the Natural Sciences and Engineering Research Council (NSERC) of Canada, the York Research Chairs (YRC) program, and Compute Canada.
Ethical Considerations
The models used for experiments are only used as the judge to evaluate other LVLM-generated responses. Therefore, the LVLM responses do not pose any ethical concerns. Additional compensation for human evaluation is not needed since it was conducted by two authors of this paper.
Appendix A Appendix
A.1 Regarding Model and Dataset Selection
With the rapid development of numerous LVLMs recently, our evaluation may not cover all LVLMs Grattafiori et al. (2024). Nonetheless, we tried to use as many state-of-the-art LVLMs as possible (we mainly focused on popular LVLMs released by the beginning of 2025, with sizes less than 10B parameters). While some LVLMs that are proposed at the time of writing this paper and not used in this study may yield different results Bai et al. (2025), the findings from our study will still offer substantial insights into building better-automated evaluators and chart-specific LLM judges in the future. Although there are many other chart benchmarks currently available, not all benchmarks (e.g., ChartQA Masry et al. (2022), fact-checking over chart images Akhtar et al. (2023, 2024)) require qualitative evaluation. Therefore we selected OpenCQA and VisText since qualitative evaluation is often required in these datasets Islam et al. (2024). Moreover, we prioritize VisText over other chart captioning datasets since it contains both L1 and L2/L3 type captions.
A.2 Prompts for the LVLM Judge
OpenCQA Pointwise (With Reference) Suppose, you are a human evaluator. You are required to rate the {Evaluation Criteria} of the answer generated by a model in comparison to the gold reference answer for a given question in the open-ended chart question answering task. Please provide your response in JSON format with the following keys: (i) Score, (ii) Explanation. Here, the ’Score’ should be between 1 to 5 (inclusive), with the higher score indicating better. Moreover, the value for the "Explanation" key will contain the reason behind your scoring. You should only provide the response in the required JSON format without any additional text. In the following, you are first given the question, followed by the gold reference answer. Afterward, you are given the model-generated answer. You are also provided with the chart image as the context for the chart question-answering task. [Question] [Gold Reference Answer] [Model Generated Answer] [Chart Image]
OpenCQA Pairwise (Without Reference) Suppose, you are a human evaluator. You are given the answers generated by two different models for a given question in the open-ended chart question answering task. Now, your task is to determine which model is better in terms of {Evaluation Criteria}. Please provide your response in JSON format with the following keys: (i) Model, (ii) Explanation, Here, the output value for the ’Model’ key is the respective model that is better, could be either ’Model A’ or ’Model B’, or ’Tie’ if both models are equally good. Moreover, the value for the "Explanation" key will contain the reason behind your preference. You should only provide the response in the required JSON format without any additional text. In the following, you are first given the question. Afterward, you are given the model-generated answers. You are also provided with the chart image as the context for the chart question-answering task. [Question] [Model 1 Generated Answer] [Model 2 Generated Answer] [Chart Image]
VisText L1 Pointwise (With Reference) Suppose, you are an human evaluator. You are required to rate the {Evaluation Criteria} of the L1 caption describing the aspects of the chart’s construction (e.g., chart type and axis labels) generated by a model in the chart captioning task. Please provide your response in JSON format with the following keys: (i) Score, (ii) Explanation. Here, the ’Score’ should be between 1 to 5 (inclusive), with the higher score indicating better. Moreover, the value for the "Explanation" key will contain the reason behind your scoring. You should only provide the response in the required JSON format without any additional text such that I can correctly parse the result from your JSON formatted response. In the following, you are first provided with the gold reference caption. Afterward, you are given the model generated caption. You are also provided with the chart image which was used as the context for the chart captioning task. [Gold Reference Caption] [Model Generated Caption] [Chart Image]
VisText L2/L3 Pairwise (No Reference) Suppose, you are a human evaluator. You are given the captions generated by two different models in the chart captioning task. Now, your task is to determine which model is better based on {Evaluation Criteria}. Please provide your response in JSON format with the following keys: (i) Model, (ii) Explanation. Here, the output value for the ’Model’ key is the respective model that is better, could be either ’Model A’ or ’Model B’, or ’Tie’ if both models are equally good. Moreover, the value for the "Explanation" key will contain the reason behind your preference. You should only provide the response in the required JSON format without any additional text such that I can correctly parse the result from your JSON formatted response. In the following, you are provided with the model generated captions. You are also provided with the chart image which was used as the context for the chart captioning task. [Model 1 Generated Caption] [Model 2 Generated Caption] [Chart Image]
A.3 Error Analysis
Some example error cases are demonstrated below.
Figure 3: An example of an error case involves the ChartGemma model being tasked with evaluating the Python code responses generated by two different models and providing a verdict on which one is better. However, instead of following the instructions, the model failed to complete the task correctly and simply returned two print statements as its output (highlighted in red text).
Figure 4: An example of an error case involves the PaliGemma-3b model being tasked with evaluating a chart caption generated by another model. Specifically, it was asked to rate the caption on a scale of 1 to 5 based on the ‘Informativeness’ criterion and to provide an explanation for the rating. However, instead of performing the evaluation correctly, the model hallucinated and repeatedly generated the same line without adhering to the required JSON format. (highlighted in red text).
Figure 5: An example of an error case for the LLaVA-Critic-7B model which demonstrates position bias by changing its selection of the better caption based on the change in the order of the model generated captions.
Figure 6: An example of an error case for the LLaVA-Critic-7B model which demonstrates length bias by selecting the lengthy caption even though in the given multidimensional prompt, conciseness was one of the criteria for a better caption.